OpenStudy (anonymous):

Modulus of complex number z= (3+4j)/(-2+5j)? Ive got to 14/29 - 23/29j then what?

OpenStudy (anonymous):

Which then becomes \[\sqrt{\left( \frac{ 14 }{ 29} \right)^{2}+\left( \frac{ 23 }{ 29 } \right)^{2}}\]

OpenStudy (anonymous):

\[\sqrt{\frac{ 725 }{ 841}}\] Is this correct, and how do I simplify further?

OpenStudy (fibonaccichick666):

I'm assuming you've simplified and multiplied correctly

OpenStudy (anonymous):

Yes pretty good at that side of things.

OpenStudy (fibonaccichick666):

Then that appears to be a good answer not further simp. now let me check it

OpenStudy (fibonaccichick666):

Check with me?

OpenStudy (fibonaccichick666):

So I'll type the tex commands, you tell me what to do ok?

OpenStudy (anonymous):

yes ok

OpenStudy (fibonaccichick666):

\[z= \frac{3+4i}{-2+5i}\]

OpenStudy (fibonaccichick666):

ok now, how do I get rid of the complex on the bottom?

OpenStudy (anonymous):

I'll show my working to where Im up to. Give me a sec.

OpenStudy (fibonaccichick666):

it's cool just work with me

OpenStudy (fibonaccichick666):

I'd prefer it ... I haven't solved it yet

OpenStudy (anonymous):

\[\[\left( \frac{( 3+4j)(-2-5j }{ (-2+5j)(-2-5j} \right)\]\] \[\frac{ -6-15j-8j-20j ^{2} }{ 4+10j-10j-25j ^{2} }\] \[\frac{ 14-23j }{ 29 }\]

OpenStudy (fibonaccichick666):

Ok, then I concur with your answer

OpenStudy (fibonaccichick666):

As for simplify, just do prime decomposition I'd say

OpenStudy (anonymous):

Its not simplified enough to be as per the multiple choice answers Im given.

OpenStudy (fibonaccichick666):

decimal form? What are your choices?